# How robust are the parameter constraints extending the $\Lambda$CDM model?

### v1

@@ Abstract
We present model-marginalized limits on the six standard $\Lambda$CDM cosmological parameters ($\Omega_{\rm c}h^{2}$, $\Omega_{\rm b}h^{2}$, $\theta_{\rm MC}$, $\tau_{\rm reio}$, $n_{s}$ and $A_{s}$), as well as on selected derived quantities ($H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$ and $r_{\rm drag}$), obtained by considering three independent Cosmic Microwave Background (CMB) experiments: the Planck satellite, the Atacama Cosmology Telescope, and South Pole Telescope. We also consider low redshift observations in the form of Baryon Acoustic Oscillation (BAO) data from the SDSS-IV eBOSS survey and Supernovae (SN) distance moduli measurements from the Pantheon-Plus catalog. The marginalized errors are stable against the different fiducial cosmologies explored in this study. The largest impact on the parameter accuracy is produced by varying the effective number of relativistic degrees of freedom ($N_{\rm eff}$) or the lensing amplitude ($A_{\rm lens}$). Nevertheless the marginalized errors on some derived parameters such as $H_{0}$ or $\Omega_{\rm m}$ can be up to two orders of magnitude larger than in the canonical $\Lambda$CDM scenario when considering only CMB data. In these cases, low redshift measurements are crucial for restoring the stability of the marginalized cosmological errors computed here. Overall, our results underscore remarkable stability in the mean values and precision of the main cosmological parameters, making irrelevant the choice of different possible cosmological scenarios once both high and low redshift probes are fully accounted for. The very same results should be understood as a tool to test exotic cosmological scenarios, as the marginalized values should be used in numerical analyses due to their robustness and slightly larger errors, providing a more realistic and conservative approach.
@@ I Introduction
As a result, one straightforward question is how stable the mean values and errors of the main cosmological parameters (including the derived ones) describing the minimal $\Lambda$CDM scenario are. Answering this question is of primary importance because robustness in their mean values and errors is absolutely required to test exotic cosmological scenarios. Therefore, a stable and robust estimation of these parameters constitutes a unique tool for cosmological analyses.
In this manuscript, we address the issue raised above by computing the marginalized errors on the main cosmological parameters over a range of possible fiducial cosmologies (see Refs. Gariazzo and Mena (2019); di Valentino et al. (2022); Di Valentino et al. (2023a, b) for previous studies dealing with such a marginalization procedure). Therefore, we not only provide an answer regarding the stability of our current (minimal) description of the universe but also offer a set of robust mean values and errors that can be used as input when testing different non-standard physical scenarios.
@@ II.1 Basic statistics
In order to perform Bayesian model comparison using the Bayes factors and evaluate the strength of preference in favor of the best model, we follow a modified version of the Jeffreys’ scale [^1] Notice that our empirical scale, summarized in Tab.1, deviates from the scale defined in the original Jeffreys’ work Jeffreys (1939)) extracted from Ref. Trotta (2008), see Tab. 1.
@@ II.2 Theoretical Models
As pointed out in the introduction, a key point in our analysis is to derive robust bounds on the basic six $\Lambda$CDM cosmological parameters and a number of derived ones ($H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$ and $r_{\rm{drag}}$) by marginalizing over a plethora of possible background cosmologies. We recall here that $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$ and $r_{\rm drag}$ is the sound horizon at the baryon drag epoch, the comoving distance a wave can travel prior to $z_{\rm drag}$, when baryons and photons decouple.
@@ III Results
Table 3: Mean values and $68\%$ CL errors on the six $\Lambda$CDM parameters as well as on some derived ones ($H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$, $r_{\rm{drag}}$ and $D_{\rm A}$) after marginalizing over a complete and large number of possible fiducial cosmologies. We illustrate the results from Planck, ACT and SPT CMB observations, either alone or combined with low redshift measurements.
Parameter Planck ACT SPT Planck+low-$z$ ACT+low-$z$ SPT+low-$z$
$\Omega_{\mathrm{b}}h^{2}$ $0.02240\pm 0.00031$ $0.02147_{-0.0010}^{+0.00071}$ $0.02220_{-0.00064}^{+0.00066}$ $0.02242_{-0.00026}^{+0.00028}$ $0.02147_{-0.00068}^{+0.00065}$ $0.02223_{-0.00061}^{+0.00066}$
$\Omega_{\mathrm{c}}h^{2}$ $0.1197_{-0.0031}^{+0.0026}$ $0.1187_{-0.016}^{+0.0059}$ $0.1171_{-0.0089}^{+0.013}$ $0.1192\pm 0.0025$ $0.1189_{-0.010}^{+0.0033}$ $0.1178_{-0.0041}^{+0.016}$
$100\theta_{\mathrm{MC}}$ $1.04095_{-0.00062}^{+0.00064}$ $1.0423_{-0.0015}^{+0.0019}$ $1.0401\pm 0.0016$ $1.04101_{-0.00059}^{+0.00058}$ $1.0422_{-0.0013}^{+0.0014}$ $1.0402_{-0.0015}^{+0.0014}$
$\tau_{\mathrm{reio}}$ $0.053_{-0.015}^{+0.016}$ $0.064\pm 0.029$ $0.061\pm 0.029$ $0.057_{-0.014}^{+0.015}$ $0.066_{-0.023}^{+0.025}$ $0.063_{-0.026}^{+0.027}$
$n_{\mathrm{s}}$ $0.9658_{-0.0089}^{+0.0093}$ $0.990_{-0.062}^{+0.033}$ $0.966_{-0.047}^{+0.048}$ $0.9670_{-0.0079}^{+0.0075}$ $0.990_{-0.039}^{+0.027}$ $0.969_{-0.033}^{+0.053}$
$\log(10^{10}A_{\mathrm{s}})$ $3.041_{-0.033}^{+0.032}$ $3.050_{-0.069}^{+0.063}$ $3.054_{-0.064}^{+0.071}$ $3.048_{-0.029}^{+0.030}$ $3.061_{-0.049}^{+0.045}$ $3.061_{-0.058}^{+0.057}$
$H_{0}$ [km/s/Mpc] $67.3_{-6.9}^{+33}$ $67_{-10}^{+33}$ $68_{-18}^{+28}$ $67.64_{-0.95}^{+1.1}$ $67.3_{-2.9}^{+1.3}$ $67.8_{-1.5}^{+4.0}$
$\Omega_{\mathrm{m}}$ $0.315_{-0.17}^{+0.077}$ $0.31_{-0.17}^{+0.11}$ $0.30_{-0.15}^{+0.33}$ $0.310\pm 0.011$ $0.312\pm 0.013$ $0.306_{-0.013}^{+0.015}$
$\sigma_{8}$ $0.810_{-0.034}^{+0.25}$ $0.830_{-0.095}^{+0.25}$ $0.80_{-0.28}^{+0.12}$ $0.811_{-0.015}^{+0.014}$ $0.832_{-0.022}^{+0.018}$ $0.811_{-0.027}^{+0.034}$
$S_{8}$ $0.832_{-0.096}^{+0.049}$ $0.848_{-0.086}^{+0.051}$ $0.79_{-0.16}^{+0.13}$ $0.827_{-0.023}^{+0.021}$ $0.847_{-0.026}^{+0.025}$ $0.821_{-0.037}^{+0.040}$
$r_{\rm drag}$ [Mpc] $147.16_{-0.78}^{+0.85}$ $148.4_{-1.6}^{+3.6}$ $147.9_{-3.3}^{+2.3}$ $147.2_{-1.9}^{+2.1}$ $148.4_{-1.3}^{+3.6}$ $147.8_{-3.8}^{+1.3}$
Table 4: As Tab. 3 but with $95\%$ CL marginalized bounds.
Tab. 3 and 4 depict the marginalized constraints (respectively at 68 and 95% CL) on the different cosmological parameters explored here arising from the different data sets. The bounds derived within each of the fiducial cosmologies described in the previous section are shown in Tabs. 5-12, which also include the minimal $\Lambda$CDM scenario.
III.1 Base $\Lambda$CDM parameters
Refer to captionRefer to caption
@@ III.1 Base $\Lambda$CDM parameters
Let us start commenting on the baryon mass-energy density, $\Omega_{\rm b}h^{2}$. Figure 1 illustrates the shift in the one-dimensional posterior probability for $\Omega_{\rm b}h^{2}$ arising from all the fiducial cosmologies and all data combinations, including also the marginalized distributions. Notice firstly that the largest departure in the uncertainty of $\Omega_{\rm b}h^{2}$ from its reference value within the $\Lambda$CDM cosmology occurs when either $N_{\rm eff}$ or $A_{\rm lens}$ are free parameters, being twice as large for the effective number of neutrino species. This was somehow expected, as the baryon mass energy density is measured from observations of the relative height of the second CMB peak with respect to the first and third peaks: if one adds baryons, the odd peaks are enhanced over the even peaks, that is, baryons make the first acoustic peak much larger than the second. The more baryons, the more the second peak is relatively suppressed, as the extra gravity provided by the baryons will enhance the compression into the potential wells. On the other hand, a larger $N_{\rm eff}$ will mostly affect the CMB power spectrum at high multipoles $\ell$, rather than at the very first peaks, i.e. at the CMB damping tail Hou et al. (2013). If $\Delta N_{\rm eff}$ increases, the Hubble parameter $H$ during radiation domination will increase as well. Baryon-photon decoupling is not an instantaneous process, leading to a diffusion damping of oscillations in the plasma. If decoupling starts at $\tau_{\rm d}$ and ends at $\tau_{\rm ls}$, during $\Delta\tau$ the radiation free streams on scale $\lambda_{\rm d}=\left(\lambda\Delta\tau\right)^{1/2}$ where $\lambda$ is the photon mean free path and $\lambda_{\rm d}$ is shorter than the thickness of the last scattering surface. As a consequence, temperature fluctuations on scales smaller than $\lambda_{d}$ are damped, because on such scales photons can spread freely both from overdensities and underdensities. The overall result is that the damping angular scale $\theta_{\rm d}=r_{\rm d}/D_{\rm A}$ is proportional to the square root of the expansion rate $\theta_{\rm d}\propto\sqrt{H}$ and consequently it increases with $\Delta N_{\rm eff}$, inducing a suppression of the peaks located at high multipoles and a smearing of the oscillations that intensifies at the CMB damping tail. Therefore, when considering $N_{\rm eff}$ as a free parameter, the bounds on the baryon energy density are loosened as they affect the CMB anisotropies in nearby multipole regions. A similar situation, albeit to a minor extent, happens with the lensing amplitude $A_{\rm lens}$. CMB temperature fluctuations get blurred due to gravitational lensing by the large scale structure of the Universe: photons from different directions are mixed and the peaks at large multipoles are smoothed. The amount of lensing is a precise prediction of the $\Lambda$CDM model: the consistency of the model can be checked by artificially increasing lensing by a factor $A_{\rm{lens}}$ Calabrese et al. (2008) (a priori an unphysical parameter). If $\Lambda$CDM consistently describes all CMB data, observations should prefer $A_{\rm{lens}}=1$. Since the effect of the lensing amplitude is also focused in the large multipoles peaks, it partly affects the extraction of $\Omega_{\rm b}h^{2}$. These larger uncertainties in $\Omega_{\rm b}h^{2}$ are observed in case of Planck, ACT and SPT data. Once low redshift measurements are accounted for, the uncertainties are reduced and are much closer to the standard $\Lambda$CDM ones. Indeed, the marginalized errors on $\Omega_{\rm b}h^{2}$ depicted in Tab. 4 are very close to the $\Lambda$CDM ones, see Tab. 5. Figure 1 shows that effect: the spread in the mean values of $\Omega_{\rm b}h^{2}$ for ACT and SPT is reduced when low-z data are added in the numerical analyses. It also shows that the mean value preferred by ACT is lower than that preferred by SPT or Planck, showing a (mild) $\sim 2\sigma$ discrepancy. Interestingly, this mismatch does not dilute when adding low-redshift observations. The reason for that is due to the higher CMB second peak amplitude preferred by ACT measurements when compared to Planck observations, implying therefore a lower baryon mass energy density with respect to that preferred by Planck data.
Refer to captionRefer to caption

### v2

@@ Abstract
We present model-marginalized limits on the six standard $\Lambda$CDM cosmological parameters ($\Omega_{\rm c}h^{2}$, $\Omega_{\rm b}h^{2}$, $\theta_{\rm MC}$, $\tau_{\rm reio}$, $n_{s}$ and $A_{s}$), as well as on selected derived quantities ($H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$ and $r_{\rm drag}$), obtained by considering several extensions of the $\Lambda$CDM model and three independent Cosmic Microwave Background (CMB) experiments: the Planck satellite, the Atacama Cosmology Telescope, and South Pole Telescope. We also consider low redshift observations in the form of Baryon Acoustic Oscillation (BAO) data from the SDSS-IV eBOSS survey and Supernovae (SN) distance moduli measurements from the Pantheon-Plus catalog. The marginalized errors are stable against the different minimal extensions of the $\Lambda$CDM model explored in this study. The largest impact on the parameter accuracy is produced by varying the effective number of relativistic degrees of freedom ($N_{\rm eff}$) or the lensing amplitude ($A_{\rm lens}$). Nevertheless, the marginalized errors on some derived parameters such as $H_{0}$ or $\Omega_{\rm m}$ can be up to two orders of magnitude larger than in the canonical $\Lambda$CDM scenario when considering only CMB data. In these cases, low redshift measurements are crucial for restoring the stability of the marginalized cosmological errors computed here. Overall, our results underscore remarkable stability in the mean values and precision of the main cosmological parameters once both high and low redshift probes are fully accounted for. The marginalized values  can be used in numerical analyses due to their robustness and slightly larger errors, providing a more realistic and conservative approach.
@@ I Introduction
As a result, one straightforward question is how stable the mean values and errors of the main cosmological parameters (including the derived ones) describing the minimal $\Lambda$CDM scenario are. Answering this question is of primary importance because these values are often regarded as reference parameters when testing exotic cosmological scenarios. Therefore, a stable and robust estimation of these parameters constitutes a unique tool for cosmological analyses.
In this manuscript, we take a first step forward to address the issue raised above by computing the marginalized errors on the main cosmological parameters over a range of possible fiducial cosmologies based on minimal extensions of the $\Lambda$CDM model (see Refs. Gariazzo and Mena (2019); di Valentino et al. (2022); Di Valentino et al. (2023a, b) for previous studies dealing with such a marginalization procedure). Before proceeding, we stress that our analysis focuses on only minimal extensions to the cosmological model. On the one hand, we acknowledge that this could be perceived as a limitation of our results because we derive constraints marginalized over models with only one (or at most two) additional degrees of freedom typically fixed in $\Lambda$CDM. Consequently, we account only for extensions that do not significantly deviate from a baseline $\Lambda$CDM cosmology. However, it would also be impractical to attempt to account for all non-standard cosmological models that deviate more substantially from $\Lambda$CDM proposed over the past few years in the literature, and we see no compelling reason to selectively include only a few of them, as this could introduce selection biases and obscure our conclusions. In this sense, our approach is also advantageous because we test if, and to what extent, the results remain robust when relaxing some of the minimal but well-motivated assumptions within $\Lambda$CDM. As a result, we not only provide an assessment of the simplest assumptions of our current (minimal) description of the universe but also offer a set of reliable mean values and errors on key cosmological parameters that, under suitable conditions, can be regarded as more realistic and conservative, serving as inputs for testing a variety of non-standard physical scenarios.
@@ II.1 Basic statistics
In order to perform Bayesian model comparison using the Bayes factors and evaluate the strength of preference in favor of the best model, we follow a modified version of the Jeffreys’ scale [^1] Notice that our empirical scale, summarized in Tab.1, deviates from the scale defined in the original Jeffreys’ work Jeffreys (1939). extracted from Ref. Trotta (2008), see Tab. 1.
@@ II.2 Theoretical Models
As pointed out in the introduction, a key point in our analysis is to derive robust bounds on the basic six $\Lambda$CDM cosmological parameters and a number of derived ones ($H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$ and $r_{\rm{drag}}$) by marginalizing over a plethora of possible background cosmologies, obtained by extending the basic 6-parameters $\Lambda$CDM scenario. We recall here that $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$ and $r_{\rm drag}$ is the sound horizon at the baryon drag epoch, the comoving distance a wave can travel prior to $z_{\rm drag}$, when baryons and photons decouple.
@@ III Results
Figure 1: Marginalized 1D and 2D posteriors involving the six standard $\Lambda$CDM parameters and other beyond-$\Lambda$CDM parameters that exhibit the most significant correlations and degeneracy lines with the former, across all the different CMB and CMB+low-$z$ datasets analyzed in this work.
@@
$(147.16_{-0.78}^{+0.85})$ $(148.4_{-1.6}^{+3.6})$ $(147.9_{-3.3}^{+2.3})$ $(147.2_{-1.9}^{+2.1})$ $(148.4_{-1.3}^{+3.6})$ $(147.8_{-3.8}^{+1.3})$
Table 3: Mean values and $68\%$ (95%) CL errors on the six $\Lambda$CDM parameters as well as on some derived ones ($H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$, $r_{\rm{drag}}$) after marginalizing over a complete and large number of possible fiducial cosmologies. We illustrate the results from Planck, ACT and SPT CMB observations, either alone or combined with low redshift measurements.
Tabs. 3 depicts the marginalized constraints (at both 68 and 95% CL) on the different cosmological parameters explored here arising from the different data sets. The bounds derived within each of the fiducial cosmologies described in the previous section are shown in Tabs. 4-11, which also include the minimal $\Lambda$CDM scenario. In what follows, for each of the six main cosmological parameters, as well as other key derived quantities, we discuss the various correlations observed across different extended cosmological models, highlighting the main correlation and degeneracy patterns between standard and non-standard cosmological parameters. Ultimately, we present the marginalized posterior distribution functions, where information from all models converges, leading to constraints on the main cosmological parameters that should be considered more robust than those derived under a specific cosmological model. We also detail, case by case, the most important differences between the constraints obtained in our model-marginalized framework and those derived from individual models. Finally, we highlight the most relevant applications of our marginalized bounds on a parameter-by-parameter basis.
@@
III.1 Baryon Energy Density $\Omega_{\rm b}h^{2}$
Let us start commenting on the baryon mass-energy density, $\Omega_{\rm b}h^{2}$. The subsection is organized as follows: we first discuss the main correlations between the baryon energy densities and the beyond-$\Lambda$CDM parameters. The aim of this subsection is twofold: (i) on one hand, we want to understand what types of phenomenologies can alter the results on $\Omega_{b}h^{2}$ inferred within a minimal $\Lambda$CDM cosmology, thereby providing a robust physical interpretation of the reasons behind these changes; and (ii) on the other hand, we want to identify the main sources of uncertainty surrounding the constraints on the baryon energy density obtained within the $\Lambda$CDM cosmology. Specifically, we will examine if and to what extent the constraining power on the baryon energy density depends on fixing other cosmological parameters beyond the standard six to their reference values expected within the baseline model of cosmology, and how relaxing these assumptions affects the constraints. Then, we adopt the Bayesian methodologies outlined in the previous sections to derive model-marginalized constraints on the baryon energy density for all the different combinations of datasets analyzed, incorporating information from all the minimal extended cosmologies under consideration. Finally, we comment on the overall consistency of the results and their applicability.
III.1.1 Main correlations with beyond $\Lambda$CDM parameters
When comparing the results for $\Omega_{b}h^{2}$ obtained within the different fiducial extended cosmologies across the six combinations of data analyzed, we observe that the largest increase in the uncertainty of $\Omega_{\rm b}h^{2}$ from its reference value within the $\Lambda$CDM cosmology occurs when either $N_{\rm eff}$ or $A_{\rm lens}$ is treated as a free parameter, with the uncertainty being twice as large when varying the effective number of neutrino species. For these extended models, we show the 1-dimensional posterior distribution functions and the 2-dimensional marginalized contours (for all six combinations of data) in the top two panels of Fig. 1 (with the top left panel showing the joint contours in the $\Omega_{b}h^{2}$ $A_{\rm lens}$ plane, and the top right panel showing the correlation in the $\Omega_{b}h^{2}$ $\Delta N_{\rm eff}$ plane).
@@ III.1.1 Main correlations with beyond $\Lambda$CDM parameters
The correlation between $N_{\rm eff}$ and $\Omega_{b}h^{2}$ was somehow expected, as the baryon mass energy density is measured from observations of the relative height of the second CMB peak with respect to the first and third peaks: if one adds baryons, the odd peaks are enhanced over the even peaks, that is, baryons make the first acoustic peak much larger than the second. The more baryons, the more the second peak is relatively suppressed, as the extra gravity provided by the baryons will enhance the compression into the potential wells. On the other hand, a larger $N_{\rm eff}$ will mostly affect the CMB power spectrum at high multipoles $\ell$, rather than at the very first peaks, i.e. at the CMB damping tail Hou et al. (2013). If $\Delta N_{\rm eff}$ increases, the Hubble parameter $H(z)$ during radiation domination will increase as well. Baryon-photon decoupling is not an instantaneous process, leading to a diffusion damping of oscillations in the plasma. If decoupling starts at $\tau_{\rm d}$ and ends at $\tau_{\rm ls}$, during $\Delta\tau$ the radiation free streams on scale $\lambda_{\rm d}=\left(\lambda\Delta\tau\right)^{1/2}$ where $\lambda$ is the photon mean free path and $\lambda_{\rm d}$ is shorter than the thickness of the last scattering surface. As a consequence, temperature fluctuations on scales smaller than $\lambda_{d}$ are damped, because on such scales photons can spread freely both from overdensities and underdensities. The overall result is that the damping angular scale $\theta_{\rm d}=r_{\rm d}/D_{\rm A}$ is proportional to the square root of the expansion rate $\theta_{\rm d}\propto\sqrt{H}$ and consequently it increases with $\Delta N_{\rm eff}$, inducing a suppression of the peaks located at high multipoles and a smearing of the oscillations that intensifies at the CMB damping tail. Therefore, when considering $N_{\rm eff}$ as a free parameter, the bounds on the baryon energy density are loosened as they affect the CMB anisotropies in nearby multipole regions.
As seen in Fig. 1, a similar situation, albeit to a minor extent, happens with the lensing amplitude $A_{\rm lens}$. CMB temperature fluctuations get blurred due to gravitational lensing by the large scale structure of the Universe: photons from different directions are mixed and the peaks at large multipoles are smoothed. The amount of lensing is a precise prediction of the $\Lambda$CDM model: the consistency of the model can be checked by artificially increasing lensing by a factor $A_{\rm{lens}}$ Calabrese et al. (2008) (a priori an unphysical parameter). If $\Lambda$CDM consistently describes all CMB data, observations should prefer $A_{\rm{lens}}=1$. Since the effect of the lensing amplitude is also focused in the large multipoles peaks, it partly affects the extraction of $\Omega_{\rm b}h^{2}$. These larger uncertainties in $\Omega_{\rm b}h^{2}$ are observed in case of Planck, ACT and SPT data. Once low redshift measurements are accounted for, the uncertainties are reduced and are much closer to the standard $\Lambda$CDM ones.
III.1.2 Marginalized Constraints
In Figure 2, we show the displacement in the one-dimensional posterior probability for $\Omega_{\rm b}h^{2}$ arising from assuming all the different fiducial cosmologies and all data combinations analyzed in this study. In the very same figure, we highlight in bold the marginalized distribution resulting from marginalizing across the different models. The results confirm the effects described in the previous subsection: specifically, the spread in the mean values of $\Omega_{\rm b}h^{2}$ for ACT and SPT is reduced when low-$z$ data are added in the numerical analyses. It also shows that the mean value preferred by ACT is lower than that preferred by SPT or Planck, indicating a (mild) $\sim 2\sigma$ discrepancy. Interestingly, this mismatch does not diminish when adding low-redshift observations. This is due to the higher CMB second peak amplitude preferred by ACT measurements compared to Planck observations, implying a lower baryon mass-energy density relative to that preferred by Planck data. We also note that the inclusion of low-redshift measurements does not lead to a significant reduction in the constraining power; for all the data combinations involving the three different CMB experiments, the constraints remain of the same order of magnitude. That said, we observe that the constraints based on Planck CMB temperature and polarization measurements remain a factor of two more constraining compared to those based on ACT and SPT. Overall, the marginalized constraints on $\Omega_{\rm b}h^{2}$ can be regarded as more robust compared to those derived within the baseline $\Lambda$CDM cosmology, as they account for marginalization over a wide range of cosmological models, many of which recast the presence of new physics prior to recombination, such as the value of $\Delta N_{\rm eff}$. Therefore, they can serve as a resilient reference value, for instance in studies aimed at testing late-time cosmologies. Typically, in these cases, a prior on $\Omega_{\rm b}h^{2}$ is adopted, either tracking the results obtained by Planck under the assumption of a $\Lambda$CDM cosmology or – to reduce model dependence – from the BBN prediction (which, however, increases the uncertainties compared to the CMB-based constraints). Since our results account for many $\Lambda$CDM extensions and possible effects, they can be considered more robust than those derived under the assumption of the minimal $\Lambda$CDM scenario and at least competitive in terms of model dependence with the prior resulting from BBN. Therefore, they can be used in those kinds of analyses.
III.2 Cold Dark Matter Energy Density $\Omega_{\mathrm{c}}h^{2}$